OpenStudy (anonymous):

The result of 6 subtracted from a number n is at least 2. What numbers are solutions? n – 2 > 6; n > 8 n – 6 ≥ 2; n ≥ 8 n + 6 ≥ 2; n ≤ 4 n + 6 ≥ 2; n ≥ 4

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

n – 6 ≥ 2; n ≥ 8

OpenStudy (anonymous):

ok thank you so much if you can i have one last question but you dont have to answer it

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

sorry im lagging really bad Which of the following equations has an infinite number of solutions? 3x – 3 = –4x 2y + 4 – y = 16 7x + 5 = 4x + 5 + 3x 6y – 2 = 2(y – 1)

OpenStudy (anonymous):

@begzat its posted

OpenStudy (anonymous):

7x + 5 = 4x + 5 + 3x if the equation winds up with an equality and no variables, then you are dealing with an infinite number of solutions this is a rule

OpenStudy (anonymous):

oh my gosh thank you so much your a lifesaver

OpenStudy (anonymous):

:D